OpenStudy (anonymous):

m=3;(-7,-4) write the eqauation for the line with the given slope and passes through the given point.

OpenStudy (anonymous):

y+4=3(x+7)

OpenStudy (anonymous):

wats the next step kinda confused on this one.

OpenStudy (anonymous):

Use the formula : y-y1=m(x-x1) (-7,-4)=(x1,y1) ,m=3

OpenStudy (anonymous):

Use the formula y-y1 = m*(x-x1) Put the values in places of slope and x1,y1 values and find the resulting equation.